The asking price for the property is $1,000,000; rents are estimated at $200,000 during the first year and are expected to grow at 5% per year. Vacancies and collection losses are expected to be 10% of rents. Operating expenses will be 35% of effective gross income. Capital expenditures will be 5% of effective gross income. A 30-year fixed rate loan fro 70 percent of the purchase price can be obtained at 10% interest rate. The property is expected to appreciate at 3% per year and is expected to be owned for seven years and then sold. The sale cost is 6% of the sale price.

The investor tells you he would also like to know how tax considerations affect your investment analysis. You determine that the building represent 90% of value and would be depreciated over 27.5 years. The potential investor indicates that he is in the 28% tax bracket. Capital gains tax rate is 20% while depreciation recapture tax rate is 25%.

A) Write down the cash flows pro forma for year 1 to 7

B) What is the investors expected before-tax internal rate of return on equity invested? Show work.

C) What is the investors expected after-tax internal rate of return on equity invested (ATIRR)? Show work.

Use financial calculator
Answer #1

(A) Cash Flow pro forma

Years 1 2 3 4 5 6 7
Rental Income 200,000 200,000 200,000 200,000 200,000 200,000 200,000
Increment (1+5%)^0 (1+5%)^1 (1+5%)^2 (1+5%)^3 (1+5%)^4 (1+5%)^5 (1+5%)^6
Total Rental Income (A) 200,000 210,000 220,500 231,525 243,101 255,256 268,019
Vacancies and Collection losses (10%)* (A) (20,000) (21,000) (22,050) (23,153) (24,310) (25,526) (26,802)
Gross Income (B) 180,000 189,000 198,450 208,372 218,791 229,730 241,217
Operating expenses (35%) (C) (63,000) (66,150) (69,458) (72,930) (76,577) (80,406) (84,426)
Interest Expense (70%*10%* $ 1 Mn) (D) (70,000) (70,000) (70,000) (70,000) (70,000) (70,000) (70,000)
Profit before tax and Dep (E=B-C-D) 47,000 52,850 58,992 65,442 72,214 79,324 86,791
Net of Tax (1-28%)*E 33,840 38,052 42,474 47,118 51,994 57,113 62,490
Savings on account of dep (F) (WN#1) 8,182 8,182 8,182 8,182 8,182 8,182 8,182
Cash flow from Operating activities (G = E+F) 42,022 46,234 50,656 55,300 60,176 65,295 70,672
Capex (5%) (9,000) (9,450) (9,923) (10,419) (10,940) (11,487) (12,061)
Sale proceeds (WN#2) 1,156,081
Capital Gain Tax (WN#3) (77,034)
Net Cash Flows 33,022 36,784 40,733 44,881 49,236 53,808 1,137,658

WN#1 Calculation of Savings on account of Depreciation Expense;

Purchase Value of Asset = 1,000,000

Building Block in above (90%) = 900,000

Life of asset = 27.5 Years

Annual Depreciation = 900,000/27.5 = 32,727

Savings (@25% recapture tax rate) = 32,727*25% = $ 8,182

Note - Depreciation on Capex is ignored.

WN#2 Calculation of Net Proceeds from Sale of Asset

Particulars Value
Purchase Cost ($) 1,000,000
Annual Increment 3% p.a.
Holding Period 7 years
Sale Consideration net of cost => (1,000,000*(1+3%)^7)*(1-6%) 1,156,081

Note - Capex is ignored.

WN#3 Capital Gains

Particulars Value
Purchase Price 1,000,000
Depreciation for 7 Years [32,727*7] (WN#1) (229,089)
Book Value at the end of 7 years  (A) 770,911
Sale Consideration net of Cost (WN#2) (B) 1,156,081
Capital Gains (B-A) 385,170
Capital Gain Tax (20%) 77,034

Note - Capex is ignored.

(B)Calculating Equity IRR (before-tax) using a financial calculator.

Years 0 1 2 3 4 5 6 7
Profit before tax and Dep (E) (Part A) 47,000 52,850 58,992 65,442 72,214 79,324 86,791
Capex (9,000) (9,450) (9,923) (10,419) (10,940) (11,487) (12,061)
Sale Proceeds 1,156,081
Net Cash Inflow before tax 38,000 43,400 49,069 55,023 61,274 67,837 1,230,811
Equity Cash Flow (300,000)
IRR 32%

(C) Calculating Equity IRR (after-tax) using a financial calculator.

Years 0 1 2 3 4 5 6 7
Equity Cash Flow (300,000)
Net Cash Flows (Part A) 33,022 36,784 40,733 44,881 49,236 53,808 1,137,658
IRR (After tax) 29%